tigerlover (tigerlover):

What is the length of each leg of the triangle below? A. B. C. D. 18 E. 13 F. 1

tigerlover (tigerlover):

\[A. \sqrt{26}\]

tigerlover (tigerlover):

\[B. 13\sqrt{2}\]

tigerlover (tigerlover):

\[C.\sqrt{13}\]